Suppose that the six-month LIBOR interest rate is 8% per annum and the three-month LIBOR interest rate is 7.5% per annum (both rates are 30/360). Estimate the three-month Eurodollar futures price quote for a contract maturing in three months.
Answer #1

The three-month Euro dollar futures price for a contract maturing in three months,

(3 * 30/360) * F(3,3) = (1 + 0.08 * 6*30/360)/(1 + 0.075 * 3 * 30/360) - 1

0.25 * F(3,3) = 1.0208588957 - 1

F(3,3) = 0.0208588957/0.25

F(3,3) = 0.0834355828

F(3,3) = 8.34355828%

Can you please upvote? Thank You :-)